OpenStudy (anonymous):

Need help . A,B,C are the angles of a triangle, then SinA + SinB - SinC = 4Sin(A/2) Sin(B/2) Cos(C/2)

OpenStudy (anonymous):

what is the question...

OpenStudy (anonymous):

prove it . thats all

OpenStudy (anonymous):

C=180-A-B sinA+sinB-sinC = sinA+sinB-sin(180-(A+B)) =sinA+sinB-sin(A+B) =sinA+sinB-sinAcosB-sinBcosA =sinA(1-cosB)+sinB(1-cosA) =2sinAsin\(^{2}\)(B/2)+2sinBsin\(^{2}\)(A/2) =4sin(A/2)sin\(^{2}\)(B/2)cos(A/2)+4sin(B/2)sin\(^{2}\)(A/2)cos(B/2) =4sin(A/2)sin(B/2) [sin(B/2)cos(A/2)+sin(A/2)cos(B/2)] =4sin(A/2)sin(B/2)sin(A/2+B/2) =4sin(A/2)sin(B/2)sin(90-C/2) =4sin(A/2)sin(B/2)cos(C/2)

OpenStudy (anonymous):

hope that is readable